OpenStudy (kittiwitti1):

Trigonometric Identities. Questions in comments First two questions have been solved sorry

OpenStudy (kittiwitti1):

Prove that these identities are true \[\frac{1-\cos{a}}{\sin{a}}=\frac{\sin{a}}{1+\cos{a}}\] \[\sin^{4}{\theta}-\cos^{4}{\theta}=\sin^{2}{\theta}-\cos^{2}{\theta}\] \[\left(\sin{x}+cos{x}\right)^{4}=\left(1+2\sin{x}\cos{x}\right)^{2}\]

OpenStudy (kittiwitti1):

Show that each equation is NOT an identity \[\sin{2x}=2\sin{x}\] \[\sec^{2}{x}+\csc^{2}{x}=1\] *side note: I don't want a complete step by step because I want to learn. Just give me hints, or start off a few steps so I know where to go from there ^_^;

OpenStudy (anthonyym):

The first equation: try cross multiplying

OpenStudy (kittiwitti1):

Sorry, I figured out #1 and #2 @anthonyym ^_^;

OpenStudy (anthonyym):

Oops sorry

OpenStudy (kittiwitti1):

It's okay thank you! :) Could you help me with the others though? o:

OpenStudy (anthonyym):

I'll try

OpenStudy (kittiwitti1):

Alright thanks :D

OpenStudy (anthonyym):

For disproving sin(2x)=2sinx I think we can plug in a number for x?

OpenStudy (kittiwitti1):

Oh! Yeah, that could work

OpenStudy (kittiwitti1):

So something like x=30 degrees?

OpenStudy (anthonyym):

How about dividing by sin(2x) then. 1 does not equal 2

OpenStudy (kittiwitti1):

Dividing by sin2x?

OpenStudy (anthonyym):

I guess, sorry I'm really just learning this right now lol

OpenStudy (kittiwitti1):

It's okay you're trying your best :D So you still get a medal :P

OpenStudy (kittiwitti1):

I'm not sure what you meant by dividing by sin2x though D:

OpenStudy (anthonyym):

I mean sin(2x) divided by sin(2x) is 1

OpenStudy (kittiwitti1):

Ah, yeah o:

OpenStudy (kittiwitti1):

Well, I was trying out the use a number for x idea. I put x=30 degrees\[\sin{(2\times30)},2\sin{30}\]That gets me:\[\sin{60},2\sin{30}\]I know those aren't equal so yeah :P